# Nonlinear Permuted Granger Causality

### v1

@@ 9.2 Proofs
To establish an bounds on the variance, we seek both lower and upper bounds for the terms in Equation 83. The bounds for the second term can be directly obtained from the derivation in the Proof of Lemma 1.
@@
For the first term in Equation 83, we follow the strategy of the proof of Lemma 1. The matrix in the trace of Equation 86 is square and positive semidefinite, and we apply the result of Remark 1.
@@
With the expected values from Condition 6, we can simplify the form to that shown in Equation 97.
@@
Taking the limit as $T\rightarrow\infty$, we obtain an upper bound for the variance of the estimate given all individual covariance matrices for one test set with each initialization and potential realization of the data. As the number of time points gets large with a fixed $T_{k}<\infty$, the size of each training set $T_{-k}$ tends to infinity.
@@
Returning to the total variance in Equation 82, with the limit in Equation 100, we can apply the Dominated Convergence Theorem to establish the limiting value for $\text{Var}\left(\hat{\vartheta}_{m}\right)$. We also use the results in Equations 63 and 72 to establish the limiting expectation $\lim_{T\rightarrow\infty}\mathbb{E}\left[T_{k}^{-1}\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]=\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)$.
The first term in Equation 82 has an inner term of the variance of the variation parameter given the initialization specific matrix $\mathbf{S}_{m,\omega,r}$ and the realization specific matrix $\bm{\Sigma}_{m,\omega}$. The only remaining source of variation arises from labelling the training and test sets $k=1,\ldots,K$. With random assignment, and fully explained temporal dependence as in Condition 3, we treat these splits as uncorrelated draws, and the variance can move inside the summation.
@@
Given the data realized covariance matrix $\bm{\Sigma}_{m,\omega}$, each initialization $\mathbf{S}_{m,\omega,r}$ is an uncorrelated draw, and we can distribute the variance term like above. From Condition 2, we extract the limiting value.
@@
The proof of Theorem 1 follows directly from the results of Lemmas 1 and 2.
Proof of Theorem 1.
From Lemma 1, we establish that $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$. Turning to the result of Lemma 2,
@@
The proof of Theorem 2 is a direct result of Condition 2.
Proof of Theorem 2.
From Condition 2, we see that $\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\overset{\textit{i.i.d.}}{\sim}\left(\vartheta_{m},\tau_{\omega}^{2}\right)$. For an individual realization, we have shown in the proofs of Lemma 1 and 2 that
@@
as the number of observations $T\rightarrow\infty$. The variation parameter is the average of the individual realization-specific covariance matrices, and the Central Limit Theorem arises from the application of Slutsky’s Theorem. ∎
The proof of Theorem 3 begins with the empirical distribution function, and employs the result of the Glivenko-Cantelli to show convergence in distribution to a $\text{Uniform}(0,1)$ random variable.
Proof of Theorem 3.
@@
From the results of Lemmas 1 and 2, we know that as $T\rightarrow\infty$, $\hat{\vartheta}_{1}\xrightarrow{d}\mathcal{F}$, where the continuous distribution $\mathcal{F}$ has expectation $\vartheta$ and variance $\varphi^{-1}\tau_{\omega}^{2}+\left(\varphi\mathscr{R}\right)^{-1}\tau_{r}^{2}$. Under the null, each realization $\hat{\vartheta}_{m}$ is independently drawn from $\mathcal{F}$. From the Glivenko-Cantelli Theorem as $T\rightarrow\infty$,
@@
Combining these results and Theorem 1 with the continuous mapping theorem in Equation 127 and Slutsky’s Theorem,
@@
The proof of Theorem 4 follows from the result of Theorem 1.
Proof of Theorem 4.
@@
From the result of Theorem 1, when $\tau_{\omega}^{2}=0$,
@@
This implies $\hat{Q}_{M}\xrightarrow{p}Q$ in both cases presented in the statement of Theorem 4. ∎
9.2.1 Finite Sample Properties
For a finite sample, we can derive the distribution of the estimated quantity $\hat{\vartheta}_{m}$ as a sum of linear combinations of chi-square random variables.
Theorem 5.
With a finite sample containing $T$ observations, $\mathscr{R}$ random initializations, and $\varphi$ realizations in the set $\Omega_{\text{obs}}$, the estimate for the variation parameter $\hat{\vartheta}_{m}$ defined in Equation 6 follows a generalized chi-square distribution of the following form
$\hskip 36.135pt+\sum_{j=d\lfloor(i-1)/d\rfloor+1}^{i-1}\left(\phi_{m,\omega,r,k,i^{\prime}j^{\prime}}+1\right)s_{m,\omega,r,k,i^{*}j^{*}}\left(Y_{m,\omega,r,j,ij}-Z_{m,\omega,r,k,ij}\right)\bigg],$ (154)
where $X_{m,\omega,r,k,i},Y_{m,\omega,r,k,ij},Z_{m,\omega,r,k,ij}\sim\chi_{1}^{2}$ for all $\omega\in\Omega_{\text{obs}}$, $r=1,\ldots,\mathscr{R}$, $k=1,\ldots,K$, $i=1,\ldots,T_{k}d$ and $j<i$, $\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}=(\phi_{m,\omega,r,k,ij})$, $\mathbf{S}_{m,\omega,r}=(s_{m,\omega,r,ij})$, $i^{\prime}=\lceil i/d\rceil$, $j^{\prime}=\lceil j/d\rceil$, $i^{*}=i\bmod d$, and $j^{*}=j\bmod d$.
The proof of Theorem 5 follows a step by step derivation of the distribution from the known starting point of Condition 3.
Proof of Theorem 5.
We begin with the distribution of the model residuals that is known from Equation 11 and Condition 3, $\mathbf{U}_{m,\omega,r}\sim\mathcal{MN}_{T\times d}\left(\mathbf{0},\mathbf{I}_{T},\mathbf{S}_{m,\omega,r}\right)$. The predicted residuals can be expressed as a function of the model residuals, like in Equation 5, and their distribution follows from that of $\mathbf{U}_{m,\omega,r}$ as in Equation 35. Like in the proof of Lemma 1, we define
@@ 9.2.1 Finite Sample Properties
For simplicity, define $\bm{\Delta}_{m,\omega,r,k}=\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\otimes\mathbf{S}_{m,\omega,r}$ and write the equivalent form $\mathbf{r}_{m,\omega,r,k}=\bm{\Delta}_{m,\omega,r,k}^{1/2}\mathbf{v}_{m,\omega,r,k}$, where $\mathbf{v}_{m,\omega,r,k}\sim\mathcal{N}T_{k}d\left(\mathbf{0},\mathbf{I}\right)$. The covaraince matrix $\bm{\Delta}_{m,\omega,r,k}=\left(\delta_{m,\omega,r,k,ij}\right)$ is symmetric and each element in the vector $\mathbf{v}_{m,\omega,r,k}=\left(v_{m,\omega,r,k,i}\right)\sim\mathcal{N}(0,1)$.
@@
To obtain the distribution for the variation parameter, we aggregate this generalized chi-square distribution over the test sets $k=1,\ldots,K$, initializations $r=1,\ldots,\mathscr{R}$, and potential realizations $\omega\in\Omega_{\text{obs}}$. ∎

### v2

@@ 6 Discussion
To establish an bounds on the variance, we seek both lower and upper bounds for the terms in Equation 102. The bounds for the second term can be directly obtained from the derivation in the Proof of Lemma 6.2.
@@
For the first term in Equation 102, we follow the strategy of the proof of Lemma 6.2. The matrix in the trace of Equation 105 is square and positive semidefinite, and we apply the result of Remark 6.5.
@@
With the expected values from Condition Additional Conditions for Theoretical Results, we can simplify the form to that shown in Equation 116.
@@
Taking the limit as $T\rightarrow\infty$, we obtain an upper bound for the variance of the estimate given all individual covariance matrices for one test set with each featurization and potential realization of the data. As the number of time points gets large with a fixed $T_{k}<\infty$, the size of each training set $T_{-k}$ tends to infinity.
@@
Returning to the total variance in Equation 101, with the limit in Equation 119, we can apply the Dominated Convergence Theorem to establish the limiting value for $\text{Var}\left(\hat{\vartheta}_{m}\right)$. We also use the results in Equations 82 and 91 to establish the limiting expectation $\lim_{T\rightarrow\infty}\mathbb{E}\left[T_{k}^{-1}\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]=\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)$.
The first term in Equation 101 has an inner term of the variance of the variation parameter given the featurization specific matrix $\mathbf{S}_{m,\omega,r}$ and the realization specific matrix $\bm{\Sigma}_{m,\omega}$. The only remaining source of variation arises from labelling the training and test sets $k=1,\ldots,K$. With random assignment, and fully explained temporal dependence as in Condition 3, we treat these splits as uncorrelated draws, and the variance can move inside the summation.
@@
Given the data realized covariance matrix $\bm{\Sigma}_{m,\omega}$, each featurization $\mathbf{S}_{m,\omega,r}$ is an uncorrelated draw, and we can distribute the variance term like above. From Condition 3, we extract the limiting value.
@@
The proof of Theorem 3.1 follows directly from the results of Lemmas 6.2 and 6.4.
Proof 6.8 (Proof of Theorem 3.1).
From Lemma 6.2, we establish that $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$. Turning to the result of Lemma 6.4,
@@
The proof of Theorem 3.2 is a direct result of Condition 3.
Proof 6.9 (Proof of Theorem 3.2).
From Condition 3, we see that $\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\overset{\textit{i.i.d.}}{\sim}\left(\vartheta_{m},\tau_{\omega}^{2}\right)$. For an individual realization, we have shown in the proofs of Lemma 6.2 and 6.4 that
@@
as the number of observations $T\rightarrow\infty$. The variation parameter is the average of the individual realization-specific covariance matrices, and the Central Limit Theorem arises from the application of Slutsky’s Theorem.
The proof of Theorem 3.3 begins with the empirical distribution function, and employs the result of the Glivenko-Cantelli to show convergence in distribution to a $\text{Uniform}(0,1)$ random variable.
Proof 6.10 (Proof of Theorem 3.3).
@@
From the results of Lemmas 6.2 and 6.4, we know that as $T\rightarrow\infty$, $\hat{\vartheta}_{1}\xrightarrow{D}\mathcal{F}$, where the continuous distribution $\mathcal{F}$ has expectation $\vartheta$ and variance $\varphi^{-1}\tau_{\omega}^{2}+\left(\varphi\mathscr{R}\right)^{-1}\tau_{r}^{2}$. Under the null, each realization $\hat{\vartheta}_{m}$ is independently drawn from $\mathcal{F}$. From the Glivenko-Cantelli Theorem as $T\rightarrow\infty$,
@@
Combining these results and Theorem 3.1 with the continuous mapping theorem in Equation 146 and Slutsky’s Theorem,
@@
The proof of Theorem 3.4 follows from the result of Theorem 3.1.
Proof 6.11 (Proof of Theorem 3.4).
@@
From the result of Theorem 3.1, when $\tau_{\omega}^{2}=0$,
@@
This implies $\hat{Q}_{M}\xrightarrow{P}Q$ in both cases presented in the statement of Theorem 3.4.
The proof of Theorem 6.1 follows a step by step derivation of the distribution from the known starting point of Condition 3.
Proof 6.12 (Proof of Theorem 6.1).
We begin with the distribution of the model residuals that is known from Equation 22 and Condition 3, $\mathbf{U}_{m,\omega,r}\sim\mathcal{MN}_{T\times d}\left(\mathbf{0},\mathbf{I}_{T},\mathbf{S}_{m,\omega,r}\right)$. The predicted residuals can be expressed as a function of the model residuals, like in Equation 12, and their distribution follows from that of $\mathbf{U}_{m,\omega,r}$ as in Equation 54. Like in the proof of Lemma 6.2, we define
@@
For simplicity, define $\bm{\Delta}_{m,\omega,r,k}=\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\otimes\mathbf{S}_{m,\omega,r}$ and write the equivalent form $\mathbf{r}_{m,\omega,r,k}=\bm{\Delta}_{m,\omega,r,k}^{1/2}\mathbf{v}_{m,\omega,r,k}$, where $\mathbf{v}_{m,\omega,r,k}\sim\mathcal{N}_{T_{k}d}\left(\mathbf{0},\mathbf{I}\right)$. The covaraince matrix $\bm{\Delta}_{m,\omega,r,k}=\left(\delta_{m,\omega,r,k,ij}\right)$ is symmetric and each element in the vector $\mathbf{v}_{m,\omega,r,k}=\left(v_{m,\omega,r,k,i}\right)\sim\mathcal{N}(0,1)$.
@@
To obtain the distribution for the variation parameter, we aggregate this generalized chi-square distribution over the test sets $k=1,\ldots,K$, featurizations $r=1,\ldots,\mathscr{R}$, and potential realizations $\omega\in\Omega_{\text{obs}}$.